A sequence starting with 1 and is formed by doubling the previous number and then subtracting 1
Answer by richwmiller(17219)

Tn=2*(T(n-1))-1
T1=1
T2=2*1-1=1
T3=2*1-1=1
T4=2*1-1=1
T5=2*1-1=1
T6=2*1-1=1
T7=2*1-1=1
T8=2*1-1=1
Every term is 1